### v1

# Analog Spiking Neuron in CMOS 28 nm Towards Large-Scale Neuromorphic Processors

Abstract
The computational complexity of deep learning algorithms has given rise to significant speed and memory challenges for the execution hardware. In energy-limited portable devices, highly efficient processing platforms are indispensable for reproducing the prowess afforded by much bulkier processing platforms.
In this work, we present a low-power Leaky Integrate-and-Fire (LIF) neuron design fabricated in TSMC’s 28 nm CMOS technology as proof of concept to build an energy-efficient mixed-signal Neuromorphic System-on-Chip (NeuroSoC). The fabricated neuron consumes 1.61 fJ/spike and occupies an active area of 34 $\mu m^{2}$, leading to a maximum spiking frequency of 300 kHz at 250 mV power supply.
These performances are used in a software model to emulate the dynamics of a Spiking Neural Network (SNN). Employing supervised backpropagation and a surrogate gradient technique, the resulting accuracy on the MNIST dataset, using 4-bit post-training quantization stands at 82.5%. The approach underscores the potential of such ASIC implementation of quantized SNNs to deliver high-performance, energy-efficient solutions to various embedded machine-learning applications.
Index Terms:
Neuromorphic circuit, spiking analog CMOS neuron, machine learning, spiking neural networks (SNN), TinyML, EdgeAI.
I Introduction
Biology-inspired neuron cells are the fundamental computing units of neuromorphic systems [1], and their energy efficiency is closely linked to them, highlighting the importance of the neuron model choice. In 1943, McCulloch and Pitts [2] proposed the first artificial neuron model, based on binary logic principles. They were followed in 1952 by Hodgkin and Huxley who developed a continuous, conductance-based, model that more accurately described the biological neuron activity [3].
The McCulloch and Pitts model was closely linked to the digital computing paradigm with 1’s and 0’s, rendering it easier to simulate on digital computers than the differential equation-based model proposed by Hodgkin and Huxley, but the inter-neural connection weights of the model were manually set, severely limiting its learning capacity. In the sixties, Mead noticed that the differential equations governing the dynamics of biological neurons bore a close resemblance to those governing the subthreshold dynamics of CMOS transistors [4]. Therefore, a few transistors operating in the analog domain could replace thousands of transistors working collectively on a digital chip to emulate differential equations.
Artificial spiking neurons aim to emulate the characteristics of biological neurons while maximizing the computational capabilities, flexibility and robustness of the neuromorphic systems that use them. Spiking neurons differ from the more common rate-coded neurons by their spiking instead of continuous output, which leads to substantial energy saving. The mathematical models are typically implemented using CMOS transistors, with each functionality incurring costs in terms of silicon footprint and power dissipation [5]. The key performance metrics of the resulting CMOS circuits consist of several parameters such as total energy consumption, chip area and biological plausibility.
Energy consumption is estimated both in terms of static power consumption and dynamic energy per spike. Additional metrics include supply voltage, mean and maximum spiking frequency, communication support (AER, SPI, I2C), spiking and bursting patterns, runtime reconfigurability (thresholds, time constants, refractory period), spike frequency adaptation, intrinsic plasticity, load driving capability, synaptic integration with non-volatile memory (NVM) devices such as memristors, and robustness against Process Variation and Temperature (PVT) variations. Achieving an appropriate balance between these parameters for the target application is crucial for constructing efficient neuromorphic systems.
The system of differential equations in the Hodgkin and Huxley model of a biophysical neuron is difficult to simulate on a large scale. In 2003, Izhikevich proposed a more efficient model in this respect, demonstrating an array of cortical spiking and bursting behaviors [6, 7]. Early silicon implementations of the model include the works in [8, 9, 10]. More recent work includes an implementation predicated on a ring oscillator in 65 nm CMOS node [11]. The spiking and bursting modes of this neuron [11] were dictated by digital control signals, which can be supplied by an FPGA. The circuit was optimized with respect to area, energy consumption, and display resilience against PVT variations as corroborated in [12].
Another spiking model is proposed in [13]. The model is described by two nonlinear differential equations linked to exponential functions. There is a membrane potential equation that features instantaneous activation of Calcium current, coupled with an additional equation that details the slower activation of Potassium current.
In yet another work [14], two neuron implementations are explained. A bio-inspired version, equipped with leak conductance, feedback loop, and related features, is proposed alongside a streamlined version that employs fewer transistors and enables a higher spiking rate. The energy expenditure per spike was evaluated at 4fJ/spike, with the spatial footprint of the simplified version determined to be 35 $\mu$m2(65nm). Subsequently, the proposed bio-inspired version was optimized on 55nm CMOS node in [15]. The first implementation, consuming a mere 2.6 fJ/spike, exhibited Fast Spiking, while the Low Threshold Spiking (LTS) second neuron with additional components, demonstrated spike frequency adaptation.
An integrator circuit at the input can generate exponential sub-threshold behavior from a steady input current and mimic leak conductance. The current-mode low-pass filter methodologies presented in academic works include Differential Pair Integrator (DPI), Tau Cell, and Low-pass filter [16]. The DPI is the most condensed and least intricate implementation. The crucial elements of such a neuron include a positive feedback circuit that simulates sodium channel activation and potassium channel deactivation for prompt toggling, a block for spike frequency adaptation, a comparator, a block to generate spikes, and a circuit for reset and refractory period.
The adaptive exponential Integrate and Fire neuron, also known as I&F neuron, includes spike frequency adaptation. Essentially, the neuron adjusts to repeated stimuli, leading to a rise in the voltage threshold and a subsequent decrease in the neuron’s firing rate. Boahen demonstrated the feasibility of spike-frequency adaptation by linking a four-transistor ”current-mirror integrator” in negative-feedback mode to any Integrate and Fire circuit [17]. This idea was further adopted by Indeveri et al. in [18], where the implementation of AdexIF neuron was showcased. This neuron concept underwent evolution over time, with feature size reduction and the incorporation of attributes such as threshold voltage and refractory period modulation. One implementation, featured in [19], employs 22 nm FDSOI technology. The AdeXIF neuron is incorporated in the DYNAPSE chip [20], representing one of the cutting-edge neuromorphic chips.
The leaky-integrate-and-fire modeling approach offers a fair compromise between computational complexity and biological features. The LIF is the simplest model to emulate a biological neuron, as it only accounts for the leaky integration and the spiking behaviour. Although many works have been published about modeling the spiking behavior of biological neurons, the LIF model is a generally accepted trade-off between the complexity of biologically plausible neuron models such as Hodgkin-Huxley and Izhikevich models [21, 22]. The simplicity of this model makes it promising for the implementation of large-scale NeuroSoC processors in real-life applications, as is the case for the IBM TrueNorth chip [23] and Intel Loihi [24]. However, they are implemented with digital architectures.
This paper explores the modeling, behavior, and energy consumption of an analog LIF neuron implemented in CMOS 28 nm technology. In Section II, the paper introduces the architecture of the LIF model and its CMOS implementation. Section III focuses on the validation and characterization of the LIF neuron design. Section IV presents the results obtained from measurements of 20 ASIC samples. Additionally, it includes a software model to emulate the analog LIF neuron, which is utilized to train a deep SNN on the MNIST dataset. This section also evaluates the energy per prediction figure of merit. Finally, Section V concludes the research presented in this paper.
II Materials
II-A LIF Model
The Leaky Integrate-and-Fire (LIF) model is based on the observation that a biological neuron essentially integrates input currents and generates output spikes (action potentials) when the membrane voltage level crosses a predefined threshold. The neuron receives an input current $I_{\text{syn}}$ from presynaptic neurons through synapses, represented by weighted connections. $I_{\text{syn}}$ increases the electrical charge inside the cell, with the cell’s membrane acting like a leaky capacitor with respect to the outside, and a resistance in parallel setting the leakage current for a given membrane potential $V_{\text{mem}}$. The latter is set by the following differential equation:
$C_{mem}\frac{dV_{\text{mem}}(t)}{dt}=I_{\text{syn}}(t)-\frac{dV_{\text{mem}}(t)}{R_{m}}$ (1)
In practice, equation 1 is discretized to enable numerical simulation. Using a time step ($dt$) and a resting membrane potential $V_{\text{reset}}$, it becomes:
$V_{\text{mem}}(t+1)=V_{\text{mem}}(t)+\frac{-(V_{\text{mem}}(t)-V_{\text{reset}})+R_{m}I_{\text{syn}}(t)}{\tau_{m}}\cdot dt$ (2)
This equation represents a balance between two forces: the leaky term $-(V_{\text{mem}}(t)-V_{\text{reset}})$, which drives the membrane potential towards the reset potential, and the input current term $R_{m}I_{\text{syn}}(t)$, which pushes the membrane potential in the opposite direction.
When $V_{\text{mem}}(t)$ reaches the threshold potential $V_{\text{th}}$, the neuron generates an output spike and the membrane potential returns to $V_{\text{reset}}$. Finally, to account for the refractory period of biological neurons, it is prevented from generating new spikes during $t_{\text{ref}}$.
The parameters that govern the LIF neuron’s operation are critical for the accurate emulation of neuronal dynamics. A clear understanding of these parameters is necessary to allow tuning the model to reflect the behaviors observed in biological systems.
  • $V_{\text{reset}}$: the voltage to which the membrane potential ($V_{\text{mem}}$) returns after a spike is generated.
  • $V_{\text{th}}$: the membrane voltage at which the neuron fires a spike.
  • $V_{\text{spike}}$: the LIF output when it generates a spike.
  • $\tau_{m}$: the time constant of the membrane potential response to input currents. It is the product of the membrane resistance ($R_{m}$) and the membrane capacitance ($C_{mem}$).
  • $dt$: the discrete time interval to update the membrane potential of the model. Smaller values result in more accurate simulations at the cost of increased computational complexity.
II-B LIF Model Computation Sequence
Algorithm 1 below provides the pseudo-code of the LIF model’s operation. It computes the membrane potential $V_{\text{mem}}$ of a neuron over time, integrating incoming synaptic currents ($I_{\text{syn}}(t)$) and producing spikes when the potential crosses a certain threshold ($V_{\text{th}}$). Initially, $V_{\text{mem}}$ is set to a reset value $V_{\text{reset}}$. Then, the model updates $V_{\text{mem}}$ at each time step ($dt$) based on the synaptic input ($I_{\text{syn}}(t)$), the membrane time constant ($\tau_{m}$) and the membrane resistance ($R_{m}$). If $V_{\text{mem}}$ exceeds $V_{\text{th}}$, it is reset to $V_{\text{reset}}$, simulating the firing of a spike, and $V_{\text{spike}}$ is output. The process incorporates a refractory period ($t_{\text{ref}}$), during which the neuron cannot fire again. This model captures the essential dynamics of neuronal spiking behavior using mathematical and logical constructs.
Algorithm 1 Leaky Integrate-and-Fire (LIF) Model
$V_{\text{reset}}$, $V_{\text{th}}$, $V_{\text{spike}}$, $\tau_{m}$, $R_{m}$, $dt$, $t_{\text{ref}}$
$V_{\text{mem}}\leftarrow V_{\text{reset}}$
for each time step $dt$ do
     Calculate $I_{\text{syn}}(t)$ as a sum of input currents from presynaptic neurons
     if $(t-t_{\text{last\_spike}})\geq t_{\text{ref}}$ then
         $V_{\text{mem}}(t+1)\leftarrow V_{\text{mem}}(t)+dt*(-(V_{\text{mem}}(t)-V_{\text{reset}})+R_{m}*I_{\text{syn}}(t))/\tau_{m}$
         if $V_{\text{mem}}(t+1)\geq V_{\text{th}}$ then
              $V_{\text{mem}}(t+1)\leftarrow V_{\text{reset}}$
              Output spike = $V_{\text{spike}}$
              $t_{\text{last\_spike}}\leftarrow t$
         else
              Output spike = 0
         end if
     end if
end for
II-C LIF Implementation in CMOS
Fig 1 provides the schematic of the LIF neuron circuit in CMOS, showing the Integration block ($M_{1}$-$M_{2}$-$C_{mem}$), the Reset block ($M_{3}$-$M_{4}$-$C_{res}$), and the Spike generation block ($M_{5}$-$M_{8}$). Each blocks is discussed next, focusing on the function of each transistor and the relationship with the mathematical model of the LIF.
Figure 1: Proposed neuron circuit in CMOS. $M_{1}$-$M_{2}$-$C_{mem}$: Integration block; $M_{3}$-$M_{4}$-$C_{res}$: Reset block; $M_{5}$-$M_{8}$: Spike generation block.
Integration Block
The Integration block consists of a current mirror made of transistors $M_{1}$ and $M_{2}$, and capacitor $C_{mem}$. The input current $I_{syn}$, which represents the neurotransmitter-induced current at the synapse between two neurons, flows through the drain of $M_{1}$ and a correlated current $I_{D}$ is generated at the drain of $M_{2}$. This current charges membrane capacitor $C_{mem}$ to develop a voltage $V_{mem}$ given by:
[…]
Which yields:
$V^{max}_{mem}(T)=\frac{I_{mem}\cdot T}{C_{mem}}+V_{reset}$ (4)
where $T$ is the integration period and $I_{mem}=I_{D}$. As the membrane voltage increases in absolute value, it approaches the threshold to trigger the spike generation block.
Spike Generation Block
The Spike generation block comprises two back-to-back CMOS inverters that generate artificial action potentials at the neuron’s output. The inverters consist of transistors $M_{5}$, $M_{6}$, $M_{7}$, and $M_{8}$. When $V_{mem}$ reaches the inversion threshold $V_{\text{inv}}$, the inverters are activated producing a spike at the output. After producing a spike, the Reset block is activated to discharge the membrane capacitor $C_{mem}$ and prepare the neuron for the next cycle. Where the inversion voltage $V_{\text{inv}}$ is expressed as [25] :
$V_{inv}=\frac{V_{DD}}{2}+\frac{1}{2}\left(V_{thn}-|V_{thp}|\right)+n_{n}V_{T}\ln\left[\sqrt{\frac{I_{0p}}{I_{0n}}}\right]$ (5)
Reset Block
The Reset block consists of transistors $M_{3}$ and $M_{4}$, and capacitor $C_{res}$. After a spike generation, this block is responsible for discharging membrane capacitor $C_{mem}$ and resetting membrane voltage $V_{mem}$ to its resting value $V_{reset}$. During the spike generation, $V_{out}$ increases across $C_{res}$ until it turns transistor $M_{4}$ on, thus discharging $C_{mem}$ and causing $V_{mem}$ to return to its resting potential, and $V_{out}$ to return to a low State. Then, the neuron is ready to start over the cycle of integration, comparison, and spike generation.
II-C1 Sub-Threshold Operation
Drain Current
Sub-threshold operation in MOS transistors occurs when the gate-to-source voltage ($V_{GS}$) is below the threshold voltage ($V_{th}$), resulting in an exponential but weak increase in the drain current ($I_{D}$) with the slight change in $V_{GS}$ [26]. Hence, the drain current requires careful analysis to understand its impact on circuit behavior. In this particular mode, the drain current of a MOS transistor is given by [25]:
$I_{D}=(n-1)I_{0}e^{\frac{q(V_{GS}-V_{th})}{nk_{B}T}}\left(1-e^{-\frac{qV_{DS}}{k_{B}T}}\right)$ (6)
Where:
$n=1+\frac{C_{\Sigma}}{C_{ox}}$ (7)
$C_{\Sigma}$ being the depletion region capacitance and $I_{0}$ given by:
$I_{0}=\mu_{0}C_{ox}\frac{W}{L}\left(n-1\right)V_{th}^{2}$ (8)
As we typically have $\left(1-e^{-\frac{qV_{DS}}{k_{B}T}}\right)\approx 1$, equation 6 can be approximated by:
$I_{D}=\mu_{0}C_{ox}\frac{W}{L}\left(\frac{C_{\Sigma}}{C_{ox}}\right)^{2}V_{th}^{2}\;e^{\left(\frac{q(V_{GS}-V_{th})}{nk_{B}T}\right)}$ (9)
Eq. (9) showcases the exponential dependency of the drain current to the gate-source voltage which is highly counter-intuitive.
Let’s apply the latter equation to the input currents $I_{syn}$ and $I_{mem}$ in Fig. 1. The source current can be modeled as a resistor since our input characterization signals are DC. Thus, $V_{SG1}=V_{DD}-R_{syn}I_{syn}$ where $R_{syn}$ is the equivalent impedance seen from the drain of $M_{1}$.
and $I_{mem}$ is given by:
$I_{\text{mem}}=I_{0P}\;e^{\left(\frac{V_{SG2}-V_{th}}{nV_{T}}\right)}$ (10)
On the other hand $V_{SG1}=V_{SG2}$:
$I_{\text{mem}}=I_{0P}\;e^{\left(\frac{V_{DD}-R_{\text{syn}}I_{\text{syn}}-V_{th}}{nV_{T}}\right)}$ (11)
Replacing in Eq. 4 while neglecting $V_{reset}$ yields:
$V^{\text{max}}_{\text{mem}}(T)=\frac{1}{C_{\text{mem}}}\cdot I_{0P}T\cdot e^{\left(\frac{V_{DD}-R_{\text{syn}}I_{\text{syn}}-V_{th}}{nV_{T}}\right)}$ (12)
Which States that $Vmem$ peak’s voltage is dependent on the input stimulation current $I_{syn}$.
Output-High Voltage of Sub-Threshold Inverter
Given the spike generation bloc of cascaded CMOS inverters $M_{5}$-$M_{6}$ and $M_{7}$-$M_{8}$, it is worth noting that the output-high voltage or $V_{OH}$ across $M_{8}$ is not necessarily equal to VDD due to the sub-threshold mode of operation [25]. We have:
$V_{OH}=V_{DD}\\ -V_{T}\ln{\left(\frac{1}{1-\frac{I_{0N}}{I_{0P}}e^{\left(\frac{-V_{DD}-V_{thn}+\mid V_{thp}\mid}{n_{n}V_{T}}\right)}}\right)}$ (13)
Where $V_{T}=\frac{k_{B}T}{q}$ is the thermal voltage approximately equal to $26mV$ at room temperature. The natural logarithmic part of the equation is negative therefore Eq. (13) yields that $V_{OH}$ is lower than $V_{DD}$. Consequently, the previous equation predicts some variability in the spike’s output voltage amplitude and the variance of the up-swing triggering in the membrane voltage.
Propagation Delay
The propagation delay of inverter $M_{5}$-$M_{6}$, $t_{P}$, is defined as the average of the low-to-high and the high-to-low propagation delays, $t_{PLH}$ and $t_{PHL}$, respectively [27]. $t_{PHL}$ can be found by averaging the discharging current of $C_{out}$ from [25]:
$t_{PHL}=\frac{(C_{gs7}+C_{gs8}+C_{gd7}+C_{gd8})V_{OH}}{I_{D6}}$ (14)
Where $C_{gs}$ is the gate-source capacitance and $C_{gd}$ the gate-drain capacitance given, respectively, by the following expressions [28]:
$C_{gs}=\frac{W}{L}C_{ox}\frac{2(1+2\eta)}{3(1+\eta)^{2}}$ (15)
$C_{total}=\frac{W}{L}C_{ox}\frac{2(\eta^{2}+2\eta)}{3(1+\eta)^{2}}$ (16)
Where $\eta$ is the drain-induced barrier lowering (DIBL) coefficient which represents how deep a transistor is in sub-threshold or how much it is not in saturation. Typical values for $\eta$ are 0.0266 and 0.0351 for the NMOS and PMOS devices [25].
Given the propagation delay dependence on the drain current and the exponential dependence of the drain current on the gate-source voltage in inverter $M_{5}$-$M_{6}$, the gate-source voltage, which is in this case the membrane voltage $V_{mem}$, will have a dissimilar effect on the inverter since the slope of $V_{mem}$ is equal to $I_{mem}/C_{mem}$.
This leads to the membrane voltage triggering at different levels because $I_{D5}$ varies exponentially hence activating the $M_{3}$ reset transistor, resulting in an up-swing of the membrane voltage up to VDD [5], at different $V_{mem}$ levels depending on it’s slope steepness.
II-D ASIC implementation
Figure 2: LIF neuron layout for implementation in TSMC 28 nm CMOS.
The proposed LIF neuron design was fabricated using the TSMC 28 nm CMOS technology node. Fig. 2 illustrates the circuit layout, which occupies an area of approximately $5~{}\mu m$ by $6~{}\mu m$. The total chip implementation has a total area of $1000~{}\mu m$ by $1000~{}\mu m$ and is encapsulated in a dual inline 20 package.
To facilitate testing after fabrication, all the inputs, outputs, and power connections are linked to the ASIC’s pads with Electro-Static Discharge (ESD) protection.
II-D1 Validation and Characterization Test Bench
Fig. 3 shows the setup used to test and characterize the fabricated ASIC. The printed Circuit Board (PCB) includes voltage regulators to provide power supply to the ASIC and the interfacing circuits. The input of the analog spiking neuron circuit is tapped to a pad on the ASIC and the latter is connected to the outside via a SMA connector on the PCB. This was done to shield the input signals from external noise.
There are also two internal analog operational amplifiers inside the chip for probing the spiking neuron’s membrane signal and spiking output. Each buffer boosts its respective signal to make it observable on an external oscilloscope as Fig. 4 shows.
All the ASIC pins associated with the input, output, and power supply of the LIF circuit are connected to the benchmark equipment using SMA connectors on the PCB. The input and power supply are connected to a Keithley 6487 picoammeter and the outputs are connected to a Keithley Infinivision MSO3024A oscilloscope as depicted in Fig. 3.
Figure 3: Experimental Setup featuring the measurement equipment and the printed circuit board test bench custom designed for the chip.
Figure 4: Internal probing buffer for the output signals of the chip. Namely the membrane integration signal and the spiking signal.
III Methods For Characterization
III-A LIF Characteristics Measurement
III-A1 Generation and integration of the LIF Synaptic Current
Figure 5: Simplified illustration of the experimental setup used to study the integration of synaptic current $I_{syn}$ in membrane capacitor $C_{mem}$ of the implemented analog LIF neuron.
The incoming synaptic current is supplied by the picoammeter which generates a DC voltage between 0 V and 200 mV to the drain of $M_{1}$ as depicted in Fig. 5, thereby setting the intensity of the input current to the LIF’s membrane. The power supply $VDD$ is set to 250 mV by a LT3045 low-dropout output (LDO) linear regulator with 500 mA driving capability, 0.8µVrms ultra-low noise (10 Hz to 100 kHz) and 76 dB PSRR at 1 MHz, and an internal analog operational amplifier in buffer configuration allows the monitoring of the membrane voltage $V_{mem}$ via the oscilloscope using a low-impedance probe.
III-A2 LIF Spiking Behavior
Figure 6: Experimental setup to study the spiking behavior of the implemented LIF neuron. The output waveform of the spiking stage is shaped by the neuron’s intrinsic properties and the characteristics of the input synaptic current.
The spike generation stage is validated by applying DC voltages at the drain and source of $M_{1}$ as shown in Fig. 6, using the same set up as previously. An on-chip analog operational amplifier buffer configuration allows to observe the output voltage $V_{out}$ node with the oscilloscope.
The subsequent behavior of the leaky integrate-and-fire (LIF) neuron can be characterized by measuring the spiking frequency as a function of the input synaptic current and systematically analyzing the relationship between these values.
III-A3 Energy Consumption per Spike Event
Figure 7: Experimental setup to characterize the energy consumption by the implemented analog LIF neuron during spiking.
This was achieved by sweeping the input voltage from 0 mV to 200 mV in 10 mV steps, using a one-second step and measuring the spiking frequency $Freq$ at the neural output with the oscilloscope and the sunk current $I_{rms}$ from the circuit with a second picoammeter plugged in series with the power supply node (see Fig. 7). This process is reiterated for each input current value, after which the root-mean-square (RMS) power consumption $P_{\text{RMS}}$ and energy per spike $E_{spike}$ are determined. The root mean square (RMS) power consumption is calculated as $P_{rms}=V_{rms}\times I_{rms}$ then the energy consumption per spike is computed as $E_{spike}=\frac{P_{rms}}{Freq}$. Using the proposed formulas, we may compute for each specific input current point, the power consumption and energy usage of the neuron circuit in Watts and Joules.
III-B Software Emulation
To validate the scaling performance of the proposed LIF architecture in a neural network, a software model was developed with the PyTorch machine learning library to simulate a network based on the measured LIF characteristics. This included defining the neural activation function based on the spiking frequency in response to the input synaptic current, measured across 20 samples of the same circuit to account for variability due to process and fabrication. The results were averaged to a single spiking function, which was then fitted using the curve fit function from Scipy library [29]. This provides a mathematical model of the activation function to implement a spiking neural network dynamic in PyTorch.
A proof-of-concept SNN was built and trained with the MNIST dataset using half-precision arithmetic. Five different feed-forward networks were experimented with as shown in table II, where the first row indicates the topology used; for instance, [400-128-10] means the first layer has 400 neurons, the hidden layer has 128 neurons and the output layer has 10 neurons.
To enhance the energy efficiency, we compressed the resolution of the MNIST dataset entries by shrinking the original 28x28 pixels image size (784 points total). Using the CV2 package and averaging each pixel with its neighboring pixels, we reduced the input data size from 784 points to 400.
The training of each network was executed on a GPU using the surrogate gradient technique, which enables to training of SNN by error backpropagation despite the non-differentiability of these networks [30, 31, 32].
Our evaluation methodology also uses Post-Training Quantization (PTQ). The model is initially trained with half-precision floating-point weights and subsequently quantized to 4-bit resolution. We then incorporate the experimentally-derived activation function from chip measurements for accuracy evaluation. This approach simultaneously enhances computational efficiency and emulates the model’s performance within a simulated resource-constrained environment. Such emulation is crucial for assessing the architecture’s real-time applicability and scalability for edge AI applications
$V_{\text{reset}},V_{\text{th}},V_{\text{spike}},\tau_{m},R_{m},dt,t_{\text{refractory}},N_{\text{syn}},W_{i,j},I_{\text{ref}},N_{i}$ (number of neurons in layer $l_{i}$)
for each layer $l_{i}$ do
     for each neuron $k$ in layer $l$ do
         $V_{\text{mem}}[k,l]\leftarrow V_{\text{reset}}$
         $t_{\text{last\_spike}}[k,l]\leftarrow-\infty$triangleright Initialize last spike time
         for each time step $t$ do
              $I_{\text{synaptic}}[k,l](t)\leftarrow\sum_{i}^{N_{i}}W_{i,j}[k]\times I_{\text{ref}}$
              if $(t-t_{\text{last\_spike}}[k,l])\geq t_{\text{refractory}}$ then
                  $V_{\text{mem}}(t+1)\leftarrow V_{\text{mem}}(t)+dt\times\left(-\frac{(V_{\text{mem}}(t)-V_{\text{reset}})+R_{m}\times I_{\text{synaptic}}[k,l](t)}{\tau_{m}}\right)$
                  if $V_{\text{mem}}(t+1)\geq V_{\text{th}}$ then
                       $V_{\text{mem}}[k,l]\leftarrow V_{\text{reset}}$
                       $V_{\text{spike}}[k,l]=1$
                       $t_{\text{last\_spike}}[k,l]\leftarrow t$
                  else
                       $V_{\text{spike}}[k,l]=0$
                  end if
              end if
         end for
     end for
end for
Algorithm 2 Fully-Connected SNN Computation Flow using LIF Model
Surrogate Gradient
The spike function of an SNN, which represent the action potential of a biological neuron, is inherently discontinuous [32, 31], and it is typically modeled using the Heaviside function in computer simulations :
$H(x)=\begin{cases}1&\text{if }x>V_{mem}^{th}\\ 0&\text{otherwise}\end{cases}$ (17)
where $V_{mem}^{th}$ is the required membrane voltage for the neuron to fire. The firing pulse is obtained by resetting $V_{mem}$ after the function value changes to 1. In order to train the SNN using the popular error backpropagation algorithm, the derivative of the Heaviside function is needed. However, this derivative is the Dirac distribution, which is 0 everywhere except at the switching threshold where it tends to infinity. This prevents the use of regular gradient descent minimization as the gradient will either be zero or at saturation during computations. [33, 34, 30].
The surrogate gradient approach addresses the problem by keeping the Heaviside function during the forward pass, and computing its derivative with a continuous approximation during the backward pass. the $Softmax$ function can be used for the purpose, with a parameter controlling its steepness near the Heaviside threshold. [30].
IV Results
TABLE I: Comparison of different neuron implementations in CMOS technology.
Parameters This work Indiveri et al. 2006 [35] Joubert et al. 2012 [36] Cruz-Albrecht et al. 2012 [37] Wijekoon et al. 2008 [38] Danneville et al. 2019 [39] Sourikopoulos et al. 2017 [14] Ferreira 2019 [15] Ronchini et al. 2020 [40] Technology node 28 nm 0.35 um 65 nm 90 nm 0.35 um 65 nm 65 nm 55 nm 180 nm Supp. Volt. (V) 0.25 3.3 V 1 V 1.2 V 3.3 0.2 0.2 (-0.1,0.1) 0.2 Implementation Analog Analog Analog Analog Analog Analog Analog Analog Analog Neuron type LIF LIF LIF LIF Izhikevich A-H ML-based ML-based Izhikevich Time scale Accelerated Biological Accelerated Biological Accelerated Biological Biological Accelerated Biological Mem. Cap. (F) 3.47 fF 432 fF 500 fF - 100 fF 5 fF 4 fF 9*934 aF 35 fF Energy (J/spike) 1.61 fJ 900 pJ 2 pJ 400 fJ 9 pJ 2 fJ 4 fJ 3.6 fJ 58.5 fJ Spiking frequency 300 kHz 200 Hz 1.9 MHz 100 Hz 1 MHz 15.6 kHz 26 kHz 360 kHz - Area/neuron (µm2) 34 2,573 120 442 2800 34 35 85 472
We present here the measurement results of the implemented LIF neuron’s characteristics with regard to the integration shape, the spiking behavior and frequency, the standby power, and the energy dissipation per spike.
IV-A LIF Synaptic Current Generation
Figure 8: Co-simulation and measurement approach to infer the LIF’s input current ($I_{SYN}$). The measured values are displayed in red, and the post-layout simulated values are displayed in blue.
The primary challenges in characterizing this architecture involve accurately measuring input current $I_{IN}$ at picoampere levels. To achieve this, the methodology proposed in this paper employs a combination of measurement and post-layout simulation to determine the input current. The input current to the LIF is sourced from a voltage source provided by the picoammeter connected to node $V_{IN}$ in Fig. 8. The voltage is applied to the shorted drain and gate of transistor $M_{1}$ via a 200 ohm resistor ($R_{ESD}$) in the ESD protection circuit. In the next paragraph, $I_{IN-sim}$ is the simulated $I_{IN}$ and $I_{IN-exp}$ is the measured $I_{IN}$ using the experimental setup.
The first phase is to perform top-level post-layout simulations to determine the relationship between the tuple ($I_{IN-sim}$, $I_{SYN-sim}$) and the resulting spiking frequency denoted $F_{SPIKE-sim}$. Then, the input current $I_{IN-exp}$ is set with the Keithley picoammeter, and the resulting spiking frequency, $F_{SPIKE-exp}$, is recorded with an oscilloscope connected to output node $V_{OUT-exp}$. The comparison of the post-layout simulation results with the measured current, which are depicted in figure 9, allows to estimate the input current at $M_{1}$’s drain $I_{SYN-exp}$ for each measured spiking frequency $F_{SPIKE-exp}$.
Figure 9: Quantitative analysis of the deviation between simulated and measured curves highlights the accuracy of the co-simulation measurement approach, further confirming its validity.
IV-B Membrane Voltage Behavior over Time
Figure 10: Measured membrane voltage over time for four synaptic current values, showing the linear integration before spike emission.
Fig. 10 illustrates the membrane voltage $V_{mem}$ response to different synaptic current levels in the proposed LIF neuron. Starting from the resting potential ($V_{reset}\approx 10$ mV), the voltage across the membrane capacitance $C_{mem}$ displays a consistent sawtooth pattern regardless of the input current magnitude. This pattern is characterized by a linear increase until the membrane voltage nears the threshold of the inverter pair $M_{5}$-$M_{6}$.
As $V_{mem}$ approaches this threshold, the activation of subsequent stages of the circuit gradually begins . This process eventually triggers transistor $M_{3}$ within the reset block, causing a sharp rise in the voltage across $M_{8}$ ($V_{out}$ up-swing), as detailed in [5].
The variations in amplitude that lead to the voltage up-swing are due to the $M_{5}$-$M_{6}$ inverter operating in a sub-threshold mode [25]. In this operational mode, the drain currents of transistors $M_{5}$ and $M_{6}$ increase exponentially with the increase in membrane voltage, which can be mathematically described by $V_{mem}=V_{GS6}=V_{SG5}$, following Eq. (9).
The exponential rise in current increases the charging rate of the gate capacitance of transistor $M_{3}$. This variable charging rate results in different activation timings for $M_{3}$, which in turn rapidly up-swings the membrane voltage $V_{mem}$. Consequently, $V_{mem}$ surpasses the inverter threshold $V^{M5-M6}_{INV}$, triggering a spike at the output. Following this spike, the rest of the circuit, particularly transistor $M_{4}$, resets $V{mem}$ back to $V_{reset}$.
Fig. 11 confirms these observations, displaying a plot of peak membrane voltages against the input current. Notably, all curves in the figure show a peak voltage of at least 60 mV, exceeding the sub-threshold swing of CMOS transistors to ensure an adequate dynamic range for the following inverter stage, which involves transistors $M_{5}$ and $M_{6}$ as shown in Fig. 8.
IV-C Spiking Behavior over Time
Fig. 11 illustrates the generated voltage spikes at different input synaptic current levels. The spike signals have a width of $1\mu s$.
The figure also shows a linear increase of the spike peak with the input current level, although not projected in equation 13. The reason is that given the technology node and the small transistor ratio’s chosen, the speed of the reset block dominates the recursive link between the spike generation inverters and the reset transistors.As varying the input current level leads to different integration slopes in capacitor $C_{mem}$, different delays occur before $V_{mem}$ reaches the switching threshold of inverter pair $M_{5}$-$M_{6}$. Then, $C_{mem}$ keeps charging and, after inverter pair $M_{7}$-$M_{8}$ also switches State, $C_{res}$ starts charging as well until it voltage reaches the threshold level to reset $C_{mem}$ through transistor pair $M_{5}$-$M_{6}$. Given the faster rise of $V_{mem}$ for larger input currents, a higher value of $V_{OH}$ across transistor $M_{8}$ is reached before $C_{mem}$ reset, hence the observed variation of the peak voltage with the intensity of the input current.
Figure 11: Measured periodicity of the generated spikes for four synaptic current levels. Each curve’s color corresponds to the indicated synaptic current value.
IV-D Spiking Frequency
Fig. 12 shows the firing frequency of the implemented LIF neuron for different input synaptic current amplitudes. The neuron’s firing frequency increases with the input synaptic current, going from 10 kHz to 350 kHz in the 10 pA to 3000 pA input range, respectively. These results were collected from 20 ASIC samples randomly selected that were wire-bonded to a Dual-In-Line 20 (DIP20) package in order to test them using the PCB test bench.
Figure 12: Measured spiking frequency for input synaptic current levels from 0 pA to 3000 pA, showing the relationship between the spike generation rate and the intensity of synaptic current.
IV-E Energy Consumption per Spike Event
Figure 13: Measured energy consumption per spike as a function of the input synaptic current for 20 different chip samples.
Fig. 13 shows the energy consumption per spike of the implemented LIF neuron for different input current values. The reported values account for both static and dynamic power consumption. Starting from about 3.3 fJ/spike corresponding to an input current of $22.6pA$, the consumed energy converges down to a pseudo-plateau averaging 2 fJ/Spike. The minimum measured energy is 1.61 fJ/spike at an input current of $1500pA$
IV-F Software emulation of the characterized analog LIF
Table II presented the results of evaluating the extracted biomimetic LIF model from our implemented analog spiking neurons to learn the MNIST dataset. The table showcases the test accuracy of different models both before and after PTQ, highlighting the biomimetic model, which achieved commendable accuracy even with low-precision parameters, indicating their suitability for efficient hardware adaptations.
TABLE II: Accuracy performance of the model executed on GPU with 4-bit quantization
Model Architecture [400-128-10] [784-128-10] [784-256-10] [784-128-10] [784-128-10]
Learning rate 0.0001 0.0001 0.0001 0.001 0.01
Neurons in hidden layer(s) [128] [128] [256] [128] [128]
Number of epochs 20 20 40 20 20
Batch size 256 256 256 256 256
Trained (32-bit) (32-bit) (32-bit) (32-bit) (32-bit)
Quantized (4-bit) (4-bit) (4-bit) (4-bit) (4-bit)
Test accuracy 84.3% 80.9% 83.6% 86.6% 88.1%
PTQ accuracy 82.5% 79% 80.6% 71.2% 67.5%
Avg. Energy per Inference (pJ) 483 753 917 679 693
V Discussion
This paper presents an implementation of a spiking neuron in a 28 nm CMOS node. The proposed design has been fabricated and comprehensively characterized. Based on this detailed characterization, a model is developed using PyTorch to enable practical emulation on GPUs at a very high neuron density.
In the proposed LIF architecture, the leak component isn’t explicitly represented by a lumped model or transistors but is instead inherent through the parasitic leakage currents of transistor $M_{4}$.
The comparison in Table II of our LIF design with published spiking neuron implementations in CMOS technology shows the achievement of our objective of a candidate for a large-scale neural network with minimal silicon footprint and energy consumption. This is shown clearly in Fig. 14 which provides a scatter diagram of recent spiking neuron implementations in CMOS technology, comparing their energy consumption per spike with their active area.
Figure 14: Area vs. Energy of analog CMOS neurons color-classified by biological model. black is Moriss-Lecar, green is AdExif, blue is Axon-Hilcock, magenta is Izhikevich and indigo is LIF. This figure depicts a scatter of CMOS neuron models according to their active area versus dynamic energy consumption per spike. The ideal spot is at the origin where the energy and the silicon footprint are minimal.
The results suggest that the proposed CMOS implementation of the LIF spiking neuron is a promising avenue for neuromorphic computing. A large-scale implemenations can be emulated by successfully deriving bio-mimetic models from analog neurons. This approach provides a robust method for verifying NeuroSOC designs before mass production and highlights the potential for creating efficient and scalable neuromorphic computing systems.
The results of this paper showcase the effective implementation and characterization of the proposed LIF architecture, which accurately integrated input synaptic currents over time, resulting in a spiking behavior that depended on the synaptic current intensity in an increasing fashion. Notably, the energy consumption per spike event was found to depend in a pseudo-linear manner on the input current level. However, there are also inherent limitations of the proposed method. The most significant are the exclusion of energy costs associated with interconnections, synaptic integration, and supporting circuits such as communication protocols. These factors, along with potential delays in weight updates and data transmission challenges, might impact the overall accuracy and efficiency of large-scale neuromorphic systems. Still, the work offers a useful foundation for future research and development in the field, especially in scaling up the LIF neuron design for extensive neuromorphic implementations.
The argument could also be made that the synaptic connections and integration are not considered in the proposed emulation method. Another argument is that the synapses are the bottleneck in a fully connected configuration. On the other hand, memristor technology is developing fast and their active footprint on silicon is much smaller than CMOS neurons. However, memristor-based devices have still a long way to go to achieve the necessary precision, reliability and robustness against PVT that will enable large-scale implementation of NeuroSoCs.
Additionally, the primary focus of this paper is to propose and showcase a scalable LIF architecture designed for CMOS implementation, aimed at facilitating the development of large-scale neuromorphic chips. The integration of synapses is out of the scope of this work.
Furthermore, the energy consumption estimate for a full neural network inference is computed based on the experimental measurement per spike. To do this estimation, the pre-trained neural network is run on the test fraction of the dataset and the number of spikes produced by each neuron in every layer is determined to obtain the total number of generated spikes by the network during one inference. Then, this value is multiplied by the average measured energy per spike of 2 fJ/Spike. The recorded results are presented in Table I. This methodology offers valuable insights into emulating and forecasting the suitability of LIF neuron design for large-scale neuromorphic systems.
There are still parameters that remain to be factored in such as potential delays in weight updates, processing of input data, and the transmission of network outputs off-chip through communication protocols such as I2C or SPI, and challenges like communication errors due to data links and noise from power sources further add to the complexity of the NeuroSoCs. In this respect, comprehensive testing and large-scale trials are necessary to fully grasp how these factors might influence the accuracy and energy efficiency of a large-scale neuromorphic system.
VI Conclusion
This paper presents an advancement in neuromorphic computing through the implementation and characterization of a Leaky Integrate-and-Fire neuron in 28 nm CMOS technology. The study’s key achievement lies in a scalable, energy-efficient, and compact architecture suitable for large-scale neuromorphic systems. By successfully fabricating and characterizing the design, and emulating a neural network that uses a software model to fit written in PyTorch, the study demonstrates a practical use to build a neural network with it.
The research also underscores the potential of CMOS technology in neuromorphic computing by comparing various spiking neuron implementations, highlighting the trend toward more efficient designs. The proof-of-concept SNN, trained on the MNIST dataset, with reduced input data size and quantized weights, emphasizes the feasibility of employing biomimetic models for efficient hardware adaptations, even with low-precision parameters. Overall, the findings from this paper suggest that the proposed CMOS implementation of the LIF spiking neuron holds great promise for advancing neuromorphic computing, offering a viable pathway for creating efficient, scalable, and robust neuromorphic systems.
Acknowledgments
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.
Authors Contributions
This should be a simple paragraph before the References to thank those individuals and institutions who have supported your work on this article.